Consider the following three cash flow series: End of Year Cash Flow Series A Cash Flow Series B Cash Flow Series C $2,420 $3,420 $2,820 $2,220 $1,620 $1,020 0 $1,000 1.5X 2.0X 2.5X 3.0X 3 2Y 2Y 2Y 4 Determine the values of X and Y so that all three cash flows are equivalent at an interest rate of 16% per year compounded yearly Carry all interim calculations to 5 decimal places and then round your final answer to the nearest dollar. The tolerance is ±5.
Answer #1
The first step is to find the PV of the Series B:
EOY Cash flow PVIF at 16% PV at 16%
0 -2420 1 -2420.00
1 3420 0.86207 2948.28
2 2820 0.74316 2095.72
3 2220 0.64066 1422.26
4 1620 0.55229 894.71
5 1020 0.47611 485.64
5426.60
To be equivalent, the PV of Series A should also be 5426.60
Hence, -1000*1+0.86207*x+1.5*0.74316*x+2*0.64066*x+2.5*0.55229*x+3.0*0.47611*x = 5426.60
Solving for x, we have
PVIF at 16% Constant with x Product
0.86207 1.00 0.86207
0.74316 1.50 1.11474
0.64066 2.00 1.28132
0.55229 2.50 1.38073
0.47611 3.00 1.42834
6.06720
The above equality becomes
6.06720*x = 6426.60
So x = 6426.60/6.0672 = $         1,059 (Answer)
CHECK:
EOY Cash flow PVIF at 16% PV at 16%
0 -1000 1 -1000.00
1 1059 0.86207 912.93
2 1589 0.74316 1180.51
3 2118 0.64066 1356.91
4 2648 0.55229 1462.19
5 3177 0.47611 1512.61
5425.16
Similarly, for Series C,
EOY Constant with y PVIF at 16% Product
0 1 1 1.00
1 1 0.86207 0.86
2 1 0.74316 0.74
3 2 0.64066 1.28
4 2 0.55229 1.10
5 2 0.47611 0.95
5.94
The equality becomes
5426.60 = 5.94*y
Y = 5426.60/5.94= $                914 (Answer)
CHECK:
EOY Cash flow PVIF at 16% PV at 16%
0 914 1 914.00
1 914 0.86207 787.93
2 914 0.74316 679.25
3 1828 0.64066 1171.12
4 1828 0.55229 1009.59
5 1825 0.47611 868.91
5430.80